# Identification by non-Gaussianity in structural threshold and smooth transition vector autoregressive models

### v2

@@ 5.1 Data
We consider a monthly U.S. dataset consisting of six variables and covering the time period from 1987:4 to 2024:2. For measuring climate policy uncertainty, we make use of the CPUI data (Gavriilidis, 2021), whose availability determines the beginning of our sample period. The CPUI is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty.
@@
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking the first differences. For measuring the price level, we use the consumer price index (CPI) and the producer price index (PPI), which are detrended by taking the first differences. As a forward looking variable, the producer price index should also help to distinguish the effects of climate policy uncertainty shocks from climate policy actions (see Gavriilidis et al., 2023). Finally, as a measure of the monetary policy stance, we include an interest rate variable (RATE), which is the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods. The CPUI and EPUI data are retrieved from https://www.policyuncertainty.com, whereas IPI, CPI, and the Federal funds rate are retrieved from the Federal reserve bank of St. Louis database and the Wu and Xia (2016) shadow rate from the Federal reserve bank of Atlanta’s website.
@@
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:2. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$), economic policy uncertainty index (divided by $10$), the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$), and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the estimated transition weights of the first regime of the fitted two-regime fifth-order LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
The series of the included variables are presented in Figure 1. Both CPUI and EPUI show a generally increased level of uncertainty towards the end of the sample period, suggesting that a model accommodating shifts in the mean could be appropriate. Moreover, there is clearly variation in the volatility of the variables, so shifts in the volatility regime should also be incorporated into the model.
5.2 Logistic STVAR model identified by non-Gaussianity
@@
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks need to be specified. As transition weights, we employ the popular logistic transition weights (Anderson and Vahid, 1998) given in Equation (2.4) for a two-regime model. They facilitate smooth transitions between the regimes depending the level of the switching variable, and they also nest the TVAR model (Tsay, 1998) as a special case when the scale parameter $\gamma$ tends to infinity. To study how the effects of the CPU shock vary depending on the level of economic policy uncertainty, we specify the first lag of EPUI as the switching variable. This approach accommodates shifts in the mean and volatility depending on the level of EPUI, and hence allows the effects of the CPU shock to vary with the level of economic policy uncertainty.
We assume that the structural shocks follow independent Student’s $t$ distributions, which is a fairly standard choice for modelling fat tailed data. Moreover, as the Gaussian distribution is obtained as a special case of the $t$-distribution, with the value of the degrees-of-freedom parameter tending to infinity, the validity of the identifying Assumption 1 can be assessed based on the estimates of the degrees-of-freedom parameters. In order to maintain a reasonable number of parameters in the model and to avoid the problems of potentially imprecise estimates and overfitting, we assume that the autoregression matrices are constant across the regimes, i.e., employ the specification in Equation (2.6). Due to the time-varying intercepts and impact matrices, the model is still fairly general, allowing for shifts in the mean, volatility, and impact responses of the variables to the shocks.
Our structural two-regime logistic STVAR (LSTVAR) model can be written as
[…]
where $c$ is the location parameter, $\gamma>0$ is the scale parameter, $y_{2t-1}$ is the lagged EPUI, the structural shocks $e_{t}=(e_{1t},...,e_{dt})$ follow independent $t$-distributions with zero mean and unit variance, and $\nu=(\nu_{1},...,\nu_{d})$ contains the degrees-of-freedom parameters of the structural shocks. Finally, we select the autoregressive order $p=5$ for our model based on AIC.
@@
The estimates of the transition weights of Regime 1 are presented in the bottom panel of Figure 1. They show that Regime 1 dominates in the periods of low economic policy uncertainty, whereas Regime 2 conversely dominates in the periods of high economic policy uncertainty. The location parameter estimate is $\hat{c}=9.80$ and the scale parameter estimate is $\hat{\gamma}=1.00$. Since the sample mean of EPUI over our sample period is $11.48$, Regime 2 starts dominating roughly when EPUI surpasses its mean, and based on the small estimate of the scale parameter, the transitions between the regimes are smooth. Therefore, we label Regime 1 as the low economic policy uncertainty (EPU) regime and Regime 2 as the high EPU regime.
@@
The estimates of the degrees of freedom parameters are small, $\hat{\nu}=(3.86,5.54,2.94,2.61,2.02,2.23)$, and therefore none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the matrices $\boldsymbol{A}_{m}$ (A.2), $m=1,2$, using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Hence, the ML estimator is expected to be well-behaved.
To assess the adequacy of our model, we utilize residual based model diagnostics to examine how well our model captures the autocorrelation structure, conditional heteroskedasticity, and distribution of the data. To that end, we plot several diagnostic figures, which are presented and discussed in more detail in Appendix B.6. They show that our model captures the autocorrelation structure of the data reasonably well, but there is a large correlation coefficient in the crosscorrelation functions between two of the residual series. Also conditional heteroskedasticity is somewhat reasonably well captured, although there are a few relatively large correlation coefficients in the auto- and crosscorrelation functions of squared standardized residuals. The marginal distributions of the series are also reasonably well captured, but the distributions of two of the residual series are somewhat asymmetric. Nonetheless, in our view, the overall adequacy of the model is reasonable enough for impulse responses analysis.
5.3 Labelling the CPU shock
The shocks of our structural LSTVAR model (5.1) are readily identified by Proposition 2. However, labelling the shocks requires external information. Moreover, due to the problem of weak identification discussed in Section 3.5, we should choose among the practically equally well fitting local solutions the one that facilitates labelling the CPU shock to the same column of the impact matrix in all of the regimes.
The estimates of $B_{1}$ and $B_{2}$ for the local solution (discussed in Section 5.2) with the highest obtained log-likelihood are the following:
[…]
where the ordering of the variables is $y_{t}=(\text{CPUI}_{t},\text{EPUI}_{t},\text{IPI}_{t},\text{CPI}_{t},\text{PPI}_{t},\text{RATE}_{t})$, and the ordering and signs of the columns of $B_{1}$ and $B_{2}$ are fixed by normalizing the entries in the first row of $\hat{B_{1}}$ to be positive and in a decreasing order.
In terms of labelling the CPU shock based on the impact responses of the variables in the estimates of $B_{1}$ and $B_{2}$, there are several properties that are reasonable to expect from it. First, since the CPU shock is the shock to CPUI, the impact response of CPUI should be substantial, and it should have the same sign in $B_{1}$ and $B_{2}$. Second, as CPU contributes to the overall economic policy uncertainty, EPUI should increase or at least not significantly decrease in response to a positive CPU shock. The effects of the CPU shock on IPI are more ambiguous. Increased uncertainty may postpone investments, decreasing production, while some firms may accelerate production to avoid future costs or invest in future technologies to gain a competitive edge. The CPU shock’s impact on the price level, in turn, may be indirect, translated through the effects on the aggregate demand, production costs, or production levels. Similarly, the Fed may not directly respond to increased uncertainty but rather to its economic impacts.
Based on the estimates of $B_{1}$ in Equation (5.2), the only variable that moves substantially at impact in response to the first shock is CPUI, and the first shock has also clearly the highest impact effect on CPUI. The first shock also moves EPUI in the same direction with CPUI, in line with increased CPU contributing to the level of general economic policy uncertainty. Also the second shock moves CPUI considerably on impact, but it moves EPUI much more, so since CPUI and EPUI vary in the same scale, it does not appear to be a plausible CPU shock candidate. Therefore, the first shock seems to be the most reasonable candidate as the CPU shock. The estimates of $B_{2}$ in Equation (5.2) show that the first shock substantially moves CPUI at impact, while the response of CPUI to the other shocks is clearly smaller. Importantly, the sign of the response of CPUI to the first shock is the same as in $\hat{B}_{1}$. Moreover, the first shock moves EPUI in the same direction with CPUI, thus, in line with what can be expected from a CPU shock. Overall, the first shock seems to be the most plausible CPU shock candidate, and hence, we deem it as the CPU shock. Since we labelled the CPUI shock in the local solution that obtained the highest likelihood, there is no need to consider other local solutions incorporating practically equal fit and mostly very similar estimates but different ordering and signs of the columns of $B_{2}$ (see Section 3.5).
@@ 5.4 Generalized impulse response functions
where $h$ is the horizon. The first term on the right side of (5.3) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the structural shock of sign and size $\delta_{i}$ in the $i$th element of $e_{t}$ hits the system at time $t$ as opposed to all shocks being random. The generalized impulse response functions also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.3).
@@
As our structural LSTVAR model has a $p$-step Markov property, conditioning on (the $\sigma$-algebra generated by) the $p$ previous observations $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ is effectively the same as conditioning on $\mathcal{F}_{t-1}$ at time $t$ and later. In order to study state dependence of the effects of the CPU shock, we obtain the histories $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$ from low EPU and high EPU regimes separately. Specifically, for each regime, we take all of the length $p$ histories in the data that imply transition weights larger than $0.75$ for the given regime, so that the given regime is clearly dominant, and the number of histories from both of the regimes is reasonably large. For low EPU regime, there are $112$ such histories and for high EPU regime there are $202$. Then, for each regime separately, we calculate the GIRFs for these length $p$ histories, and take the sample mean to obtain a point estimate for the GIRF conditional on dominance of the given regime, whereas taking sample quantiles gives confidence intervals that reflect uncertainty about the initial state of the economy within the dominating regime. This way the GIRFs estimate the effects of the CPU shock in times of low EPU and high EPU in a way that reflects the observed data. Our Monte Carlo algorithm for computing the GIRF is similar to the one described in Lanne and Virolainen (2024), Appendix B.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of Regime 1 are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to a positive one-standard-error CPU shock with initial values from Regime 1, whereas the right column shows the responses with initial values from Regime 2. All GIRFs have been scaled so that the instantaneous increase in CPUI is $5$. The shaded areas depict the $95\%$ confidence intervals that reflect uncertainty about the initial state of the economy within the regime.
The GIRFs to a one-standard-error positive CPU shock are presented in Figure 2. We also computed the GIRFs for two-standard-error and negative CPU shocks, but the responses are similar to the positive one-standard-error shock (after adjusting for the different sign or size), so they are not presented for brevity. The responses of CPUI and EPUI follow similar patterns during times of low and high EPU, with CPUI responding substantially and EPU more weakly. The responses of the macroeconomic variables, however, differ across the regimes.
When the level of EPU is low, production decreases following a positive CPU shock, possibly because of firms postponing investments (see Fried et al., 2021). On the other hand, when EPU is high, production slightly increases. A possible explanation is that some firms, already operating in a more unpredictable environment, continue with planned investments, or hedge against multiple types of uncertainties by ramping up production in sectors less likely to be immediately affected or by investing in a broader range of technologies.
Consumer prices initially decrease and then slightly increase when EPU is low, while they increase more significantly when EPU is high. The initial decrease in prices under low EPU could be due to reduced consumer demand as people postpone consumption in anticipation of new policies. Under high EPU, prices rise more substantially, possibly because of firms adapting their operations to meet potential new regulations, increasing production costs and passing these costs onto consumers.
Producer prices rise permanently in both regimes, possibly due to anticipatory adjustments to future costs. The effects of the CPU shock on the interest rate are minor but consistent with the responses of the other variables, with interest rates increasing as prices rise, more so during high EPU. Finally, the responses of the transition weights indicate a shift towards the high EPU regime, which is expected, as a positive CPU shock increases EPUI.
Overall, our results suggest that a positive CPU shock decreases production in times of low EPU but slightly increases it in times of high EPU. Prices and the interest rate increase in both regimes, but the increase in consumer prices (and interest rate) is more significant when level of EPU is high. Therefore, our results are complementary to Gavriilidis et al. (2023), who found a positive CPU shock decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.
6 Conclusions
@@
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous or, for a $p$th order model, functions of the preceding $p$ observations. Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
The identification is often weak with respect to the ordering and signs of the columns of the impact matrices in some regimes, and an incorrect ordering or signs of the columns results in impulse response functions that do not estimate the effects of the intended shock but of a linear combination of the structural shocks. In other words, there is often multiple local maxima of the log-likelihood function with almost equal log-likelihood that are very similar to each other but incorporate different ordering and signs of the columns of the impact matrices in some regimes. As a solution, we proposed a procedure that utilizes the almost equally well fitting local solutions obtainable by our estimation method. If necessary, the identification can also be strengthened by imposing economic overidentifying short-run restrictions, which are testable.
@@
In addition to discussing the identification of the shocks, we discussed the estimation of the model parameters by the method maximum likelihood in case of independent Student’s $t$ shocks, and proposed a two-phase estimation procedure that can produce a set of alternative local solutions. By making use of the results of Saikkonen (2008) and Kheifets and Saikkonen (2020), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model. While useful for obtaining a well-behaving estimator, stationarity is not, however, required for applicability of our identification results. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2024), which is available through the CRAN repository.
@@
Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S. data from 1987:4 to 2024:2. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable and allowing for smooth transitions in the intercept parameters as well as in the impact matrix. We found that a positive CPU shock decreases production in times of low economic policy uncertainty but slightly increases it in times of high economic policy uncertainty. Consumer prices, in turn, increase more significantly when the level of economic policy uncertainty is high than when it is low. Our results are, therefore, complementary to Gavriilidis et al. (2023), who found a positive CPU decreasing real activity and increasing prices, thus, transmitting to the U.S. economy as a supply shock.

### v3

@@ 5.1 Data
We consider a monthly U.S. dataset consisting of six variables and covering the time period from 1987:4 to 2024:12. Following Khalil and Strobel (2023) and Huang and Punzi (2024), we measure CPU with the climate policy uncertainty index (CPUI) of Gavriilidis (2021), which is constructed based on the amount of newspaper coverage on climate policy uncertainty topics. The availability of the CPUI data determines the beginning of our sample period. In order to control for general economic policy uncertainty, we also include the economic policy uncertainty index (EPUI) (Baker et al., 2016), which is based on newspaper coverage on topics related to economic policy uncertainty as well as on components quantifying the present value of future scheduled tax code expirations and disagreement among professional forecasters over future government purchases and consumer prices.
@@
As a measure of real economic activity, we include the log of industrial production index (IPI), which is detrended by taking first differences. For measuring the price level, we use the logs of consumer price index (CPI) and producer price index (PPI), likewise detrended by taking first differences. Finally, we measure the monetary policy stance with the effective Federal funds rate that is replaced by the Wu and Xia (2016) shadow rate for the zero-lower-bound periods (RATE). The series of the included variables are presented in Figure 1.
@@
Figure 1: Monthly U.S. time series covering the period from 1987:4 to 2024:12. From top to bottom, the variables are the climate policy uncertainty index (divided by $10$); economic policy uncertainty index (divided by $10$); the log-differences of industrial production index, consumer price index, and producer price index (multiplied by $100$); and the effective federal funds rate replaced by the Wu and Xia (2016) shadow rate for the zero lower bound periods. The bottom panel shows the transition weights of High EPU Regime (Regime 2) of the fitted two-regime second-order structural LSTVAR model. The shaded areas indicate the U.S. recessions defined by the NBER.
5.2 The structural LSTVAR model
@@
In order to specify the structural STVAR model (3.2)-(3.3), the transition weights and the distributions of the structural shocks must be specified. We employ the logistic transition weights with two regimes (Anderson and Vahid, 1998) given in Equation (2.4). They facilitate smooth transitions between the regimes depending the level of the switching variable, enabling an analysis of how the effects of the CPU shock vary depending on the level of EPU by specifying its first lag as the switching variable. The logistic weight function also has the convenient property that is nests the threshold weight function (Tsay, 1998) as a special case when the scale parameter tends to infinity, effectively incorporating selection between these two weight functions within estimation.
We estimate the structural LSTVAR model by PML with the three-phase procedure proposed in Section 4, assuming that the structural shocks follow the independent skewed $t$-distributions. The skewed $t$-distribution can flexibly capture fat tails and skewness, and as the Gaussian distribution is obtained as a special case of the skewed $t$-distribution, with $\nu_{i}=\infty$ and $\lambda_{i}=0$, the validity of the identifying Assumption 1 can be assessed based on the estimates of these parameters. To avoid overfitting, particularly in the regime prevailing when the level of EPU is high, the autoregressive order $p=2$ is selected based on Hannan-Quinn information criterion.
To deal with the problem of weak identification with respect to the ordering and signs of the columns of $B_{2}$, discussed in Section 3.3, 3.4, and 3.5, we implement the following, simplistic blended identification strategy that avoids making restrictive assumptions about the effects of the CPU shock. We normalize the first row of $B_{1}$ positive with CPUI ordered first and assume that in both regimes, one of the shocks, deemed as the CPU shock, has greater impact effect on the CPUI than any other shock, and this impact effect has the same sign in both regimes. Then, to support our identification, we check that the recovered CPU shocks align with several major historical events affecting CPU. In particular, we find that President Trump’s election is preceded by large CPU shocks in October 2016 and 2024 ($17$ and $26$, respectively). The largest negative recovered CPU shock ($-10$), in turn, dates to August 2022, when the Inflation Reduction Act was passed, providing support for the clean energy transition. Among all the shocks recovered from the model, only the one deemed as the CPU shock has signs consistent with the above historical events affecting CPU.
@@
The location parameter estimate is $\hat{c}=14.48$, implying roughly that Regime 1 dominates when the level of EPU is not high, while Regime 2 starts prevailing in times of high EPU. Therefore, we label Regime 1 as Low EPU Regime and Regime 2 as High EPU Regime. The scale parameter estimate is $\hat{\gamma}=16.65$, indicating that transitions between the regimes are somewhat fast (rather than very smooth), which can be observed also from the bottom panel of Figure 1 where the fitted transition weights of High EPU Regime are presented.
@@
The estimates of the degrees-of-freedom and skewness parameters are (2.33,2.10,2.64,3.23,6.54,(2.33,2.10,2.64,3.23,6.54,( 2.33 , 2.10 , 2.64 , 3.23 , 6.54 , 2.68)2.68)2.68 ) and $(0.41,0.17,0.09,0.43,0.25,-0.12)$, respectively. Clearly, none of the shocks is close to Gaussian, suggesting that Assumption 1 is satisfied. Then, we check whether our model satisfies the stationarity condition of Theorem A.1 in Appendix A by computing an upper bound for the joint spectral radius of the companion form AR matrices of the regimes using the JSR toolbox (Jungers, 2023) in MATLAB. The obtained upper bounds is $0.98$, which is strictly less than one, implying that our model satisfies Assumption A.1 (in Appendix A) and is ergodic stationary by Theorem A.1. Based on the residual diagnostics presented in Appendix E, the adequacy of our model to capture the autocorrelation structure, conditional heteroskedasticity, and distribution of the data seems, overall, reasonable.
@@ 5.3 Generalized impulse response functions
where $h$ is the horizon. The first term on the right side of (5.1) is the expected realization of the process at time $t+h$ conditionally on a structural shock of sign and size $\delta_{i}\in\mathbb{R}$ in the $i$th element of $e_{t}$ at time $t$ and the previous observations. The latter term on the right side is the expected realization of the process conditionally on the previous observations only. The GIRF thus expresses the expected difference in the future outcomes when the $i$th structural shock of sign and size $\delta_{i}$ arrives at time $t$ as opposed to all shocks being random. The GIRFs also facilitate tracking the effects of the shock on the transition weights $\alpha_{m,t}$, $m=1,...,M$, by replacing $y_{t+h}$ with $\alpha_{m,t+h}$ on the right side of Equation (5.1).
@@
Due to the $p$-step Markov property of our structural LSTVAR model, conditioning on the previous $p$ observations, $\boldsymbol{y}_{t-1}=(y_{t-1},...,y_{t-p})$, is equivalent to conditioning on $\mathcal{F}_{t-1}$. To the study the state-dependence of the effects of the CPU shock, we follow a procedure similar to Lanne and Virolainen (2025) and calculate the GIRFs conditional on a given regime dominating when the shock arrives. Specifically, for each Regime $\tilde{m}\in\{1,2\}$, we take all the length $p$ histories $\boldsymbol{y}_{t-1}$ from the data for which the corresponding transition weight $\alpha_{\tilde{m},t}$ is greater than $0.75$, indicating that Regime $\tilde{m}$ is clearly dominant (there are $366$ such histories for Low EPU Regime and $80$ for High EPU Regime). For each such history, we take the corresponding CPU shock recovered from the data, and compute the GIRF using the Monte Carlo algorithm described in Lanne and Virolainen (2025), Appendix B. Finally, GIRFs corresponding to shocks of different sign and size are made comparable by scaling them to correspond to a $5$ point instantaneous increase of CPUI.
@@
Figure 2: Generalized impulse response functions to the identified CPU shock $h=0,1,....,36$ months ahead. From top to bottom, the responses of CPUI, EPUI, IPI, CPI, PPI, the interest rate variable, and transition weights of High EPU Regime are depicted in each row. The responses of IPI, CPI, and PPI are accumulated to (100×100$\times$100 ×)log-levels. The left column shows the responses to conditionally on High EPU Regime dominating when the shock arrives, whereas the right column shows the responses conditionally on High EPU Regime dominating. All the GIRFs have been scaled so that the instantaneous increase in CPUI is $5$ points. Each GIRF is computed with a Monte Carlo algorithm similar to the one described in Lanne and Virolainen (2025, Appendix B).
The distribution of the GIRFs in each regime is illustrate in a so-called ”shotgun plot” in Figure 2, which depicts each GIRFs using a level of opacity such that the darkness of a region displays the concentration of GIRFs in it (cf. Lanne and Virolainen, 2025). The response of CPUI to a positive CPU shock follows a similar pattern in both regimes in the vast majority of the GIRFs, while EPUI temporarily increases in Low EPU Regime and decreases in High EPU Regime. A possible explanation for the decrease in the EPUI could be that under already heightened EPU, an unexpected increase in CPU may induce a shift in focus from broader economic policy issues to climate policy.
Consistent with the previous literature (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024), a positive CPU shock decreases production in both regimes. Huang and Punzi (2024) attribute the decrease in production to firms postponing investments, whereas Fried et al. (2022) and Khalil and Strobel (2023) consider mechanisms that shift investments toward cleaner capital. Although many of the GIRFs exhibit a short-term increase of IPI in High EPU Regime, the negative long-term response of IPI seems, on average, roughly similar in both of the regimes.
In both regimes, a positive CPU shock increases consumer prices, aligning with Huang and Punzi (2024) and Khalil and Strobel (2023). Huang and Punzi (2024) found that a positive CPU shock increases inflation by acting as a positive supply shock, whereas according Khalil and Strobel (2023), a positive CPU shock may decrease aggregate demand due to risk-averse households raising savings, while firms raise their prices pre-emptively to lower the probability of being stuck with a negative markup. However, we find that the inflationary effects are stronger in High EPU Regime than in Low EPU Regime. One possible explanation is that the firms are more risk averse in the periods of high EPU, thereby reacting more sensitively and passing the increased costs or markups onto consumers. Producer prices, in turn, mostly decrease in Low EPU Regime but increase in High EPU Regime.
The response of the interest rate variable seems consistent with the inflationary effects of the CPU shock: it increases both regimes, but clearly more so in High EPU Regime. Finally, consistent with the GIRFs of EPUI, a positive CPU shock temporarily increases the transition weights of High EPU Regime in the periods of low EPU, while the effect is opposite in the periods of high EPU. Overall, our results suggest that a positive CPU shock decreases production and increases inflation in both regimes, but its inflationary effects are stronger in the periods of high EPU than in times of low EPU.
6 Conclusion
@@
Linear structural vector autoregressive models can be identified statistically by non-Gaussianity without imposing restrictions on the model if the shocks are mutually independent and at most one of them is Gaussian (Lanne et al., 2017). We have shown that this result extends to structural threshold and smooth transition vector autoregressive models incorporating a time-varying impact matrix defined as a weighted sum of the impact matrices of the regimes. The weights of the impact matrices are the transition weights of the regimes, which we assume to be either exogenous, logistic, or of the threshold form, but our results can be generalized to other suitable weight functions as well. Thus, the types of transition weights typically used in macroeconomic applications are accommodated.
@@
In addition to discussing the identification of the shocks, we discussed the problem of labelling them, and how to utilize blended identification (cf. Carriero et al., 2024) to solve the problem of weak identification with respect to the ordering and signs of the columns of the regime-specific impact matrices. We discussed estimation of the model parameters by the method of penalized maximum likelihood, and proposed a three-phase estimation procedure by adapting the procedure of Lanne et al. (2017) to nonlinear SVAR models. By making use of the results of Saikkonen (2008), we also presented a sufficient condition for establishing stationarity, ergodicity, and mixing properties of our structural STVAR model. While useful for obtaining a well-behaving estimator, stationarity is not, however, required for the applicability of our identification results. The introduced methods have been implemented in the accompanying R package sstvars (Virolainen, 2025a), which is available through the CRAN repository.
@@
Our empirical application studied the macroeconomic effects of the climate policy uncertainty shock and considered monthly U.S. data from 1987:4 to 2024:12. As a measure of climate policy uncertainty, we employed the CPU index (Gavriilidis, 2021) constructed based on the amount of newspaper coverage on climate policy uncertainty topics. We fitted a two-regime logistic STVAR model to the data using the first lag of the economic policy uncertainty index (Baker et al., 2016) as the switching variable. We found that a positive CPU shock decreases production and increases inflation in times of both low and high EPU, but its inflationary effects are stronger in the periods of high EPU. Our results are, therefore, in line with the previous literature suggesting that a positive CPU shock decreases production (Fried et al., 2022, Khalil and Strobel, 2023, and Huang and Punzi, 2024) and increases inflation (Khalil and Strobel, 2023, and Huang and Punzi, 2024).